# Data augmentation with automated machine learning: approaches and performance comparison with classical data augmentation methods

### v2

@@ 5.6 Searchless optimization of data augmentation policies
The optimization techniques discussed in Subsections 5.1 through 5.4 – though have demonstrated impressive results – are still too expensive for most practical purposes. For instance, model optimization methods employed in AutoAugment takes about 15,000 GPU hours to train on ImageNet [125]. To address the challenge of computational complexity, some recent works (e.g., [31, 167, 168, 159]) propose simplified search space design where good augmentation policies can be found using much simpler search strategies like random search [169], grid search [170] or greedy search [171]. Greedy AutoAugment [167] proposes to reduce the search space by eliminating the need to search for different combinations of augmentation operations in a combinatorial manner. The basic idea is to dynamically expand the space of augmentation operations in the direction of effective augmentation policies by creating and modifying sub-policies with good performance. In the search process, only transformation type and intensity are the only hyperparameteres considered. The probability of applying the operations are determined later when the search is complete and is only applied on the effective policies found. This significantly reduces the amount of computations needed to find good augmentation strategies. Because of its simplicity and effectiveness, the greedy search policy optimization procedure has been used in several works [172].
@@
Unlike in RA [31] and all previous optimization strategy formulations, Uniform Augment (UA) [159] completely avoids the need to incorporate a search stage and instead propose to uniformly sample augmentation policy hyperparameters from the search space. To achieve this, the authors formulated the search space as distribution invariant by ensuring that the transformation operations and associated intensities results in label-preserving augmentations, and the transformed data remains within the distribution of the original input data. A machine learning model is then able to learn optimal augmentations by sampling randomly from the search space and optimizing model parameters using gradient descent. Uniform Augment, thus far, is the cheapest in terms of the number of computations needed to find optimal augmentation policies. Indeed, the number of search operations is theoretically zero.
@@ 6 Evaluation methods
An important step after selecting a subset of operations in the search space and an optimization method is to evaluate the performance of the resulting components. This process is repeated for different combinations of model configuration settings (i.e., all plausible sets of augmentation operations, hyperparameters and search methods). At the end of the process, performance results are compared and an optimal pipeline and hyperparameters chosen according to a specified performance criterion. This process is and takes enormous time and computational resources to achieve optimal results. The costly nature of the evaluation stage significantly limits the range of plausible configurations that can be explored to achieve optimal predictive performance. Instead of training the network to converge before evaluating its performance, approaches have been devised to accelerate the process.
One of the most popular approaches is the so-called early stopping strategy [20], a technique based on terminating the evaluation process for settings that are predicted to perform relatively poor on a validation set. Some works, for instance, RandAugment (RA) [31], TrivialAugment (TA) [168] and OHLAA [111], reduce the range of augmentation operations and hyperparameters in the search space in the evaluation process. Another common approach to speed up model evaluation is to reduce the fidelity of the input data [20]. In image augmentation, this is achieved by reducing the sizes or resolutions of the images used. Another common approach is to train on proxy tasks–i.e., employ reduced model size or subsets of training data. Examples of AutoML-based data augmentation models that employ this strategy include AutoAugment (AA) [17], FastAA [105], [105], FasterAA [27], DADA [156], AWS [142], and PBA [34]. Another method for reducing the enormous computaional overhead at the evaluation stage is to employ surrogate models as evaluators [173, 174] to predict the performance of the target models. This approach circumvents the need to perform costly evaluation on different configurations of the real model.
@@ 7 Quantitative performance of data augmentation approaches
We present results on the performance of AutoML-based data augmentation methods and compare these with the state-of-the data augmentation methods based on classical approaches. For fairness we compare methods that use the same datasets and similar training settings (i.e., similar model configuration and number of training epochs) The results are curated from the original works. We highlight instances where results or other performance information are taken from secondary sources are used.
@@ 7.1.1 Datasets
The CIFAR-10 [108] and CIFAR-100 [108] are another example of popular image classification datasets. Each of the datasets contains 60,000 labeled 32x32 RGB images. The data is divided into 50,000 training and 10,000 test images. The content of each dataset is basically generic objects The images in CIFAR-10 are divided into 10 classes of 6,000 images per class, while CIFAR-100 has 100 classes, each containing 600 (500 training and 10 test) images.
The Street View House Numbers (SVHN) [109] is a digit (i.e., decimal numbers 0 t0 9) recognition dataset containing 600,000 color (RGB) images. The dataset is made up of cropped, 32×32 images derived from Google Street View [177]. Thus, the images are pictures of real-world house number plates in challenging settings. SVHN include bounding box information in a separate file and can therefore be used for digit detection in addition to image classification.
@@
The Microsoft Common Objects in Context (MS COCO) dataset [176] large-scale dataset which can be used for many computer vision tasks. It is primarily designed for object detection, pose estimation, keypoint detection, image captioning, and segmentation tasks. It has a total of 330,000 images split into 80 classes.
@@ 7.1.2 Performance metrics and model settings
For the classification tasks, the performance metric used is the generalization accuracy. Results for ImageNet include top-1% and top-5% accuracy measures. Mean average precision (mAP) is used to characterize performance of the object detection tasks, i.e., on the MS COCO dataset (Table V).
Performance on CIFAR-10 and CIFAR-100 have been tested using three different backbone models: Wide Residual Networks [179], specifically, Wide-ResNet-28-10 (see the original work for a detailed description of this setting); Shake-Shake (26 2x32d) [180]; and PyramidNet [181]. Results for ImageNet dataset have are based on ResNet-50 and ResNet-200 backbone models (see He et al. [182] for details). RetinaNet [183] is used with various ResNet models [184] on the MS COCO dataset.
@@ 7.2 Performance of AutoML-based data augmentation methods
Most of the existing works report results for image classification using large-scale CNN models. The most popular datasets used for these purposes are the ImageNet and CIFAR (CIFAR-10 and CIFAR-100) family of datasets described in Subsection 7.1. Detailed performance results for automated data augmentation methods based on these datasets are summarized in Tables Tables III, IV and VII. Table III show image classification results for CIFAR-10 and CIFAR-10 for three different backbone models– WideResNet 28-10 (WRN28-10) [179], Shake-Shake 26-32 (S-S 26-32) [180] and PyramidNet (PNet) [181]. For each data augmentation method, we indicate the average classification accuracy for the three backbone models. From the Table, the mean gain in accuracy range between 0.6 and 1.5 on CIFAR-10 and between 0.8% and 3.1% on CIFAR-100. This demonstrates consistentently strong performance for all methods. Indeed, the average performance gain for all the methods is 1.02% and 2.17% on CIFAR-10 and CIFAR-100, respectively. Note that for CIFAR-100, many of the methods do not have results for PyramidNet backbone on CIFAR-100. We exclude these methods in our average performance computations. Similar to Table III, Table IV shows classification results for the ImageNet dataset based on ResNet-50 and ResNet-200 backbone models. Results are shown for top-1 and top-5% accuracies. The table also indicates percentage performance gain for each of these accuracy measures. The gains are consistent for all methods, with the lowest top-1% accuracy on RestNet-50 being +0.9% (for PBA method) and +1.5% on ResNet-200. The performance gains for various AutoML-based data augmentation methods on the SVHN dataset are summarized in Figure 14. Again, the results show impressive performance on challenging digit classification tasks. In Table VII we compute the average performance of different categories of automated data augmentation methods based on the optimization strategies they employ. The results show that performance difference across different search methods, on average, is largely insignificant.
While the majority of approaches have reported results for image classification tasks, a few works have evaluated the performance of automated data augmentation techniques on machine learning tasks other than image classification. For instance, Liu et al. [104], Cubuk et al. [31] and Chen et al. [185] investigate the performance of common automated data augmentation methods–AA [17], DADA [28], DDAS [104]–on object detection tasks. The tests were conducted using large-scale detection models, specifically, ResNet-101 and RetinaNet. The results of these experiments are summarized in Table V. These studies show that AutoML-based data augmentation methods are effective for object detection. Indeed, per the results in Table V, the data augmentation techniques results in an average performance improvement of 1.54% (mean average precision or mAP) for object detection models.
@@ 7.3.2 Comparison of automated data augmentation methods and state-of-the art classical approaches
In this section, we compare the performance of state-of-the-art data agumentation methods based on classical approaches with automated data augmentation methods. The results are summarized in Figures 15, 16 and 17. Again, to ensure a uniform point of reference for the assessment, we selected the same datasets and CNN backbone models for comparison. Specifically, the results are provided for CIFAR-10 and CIFAR-100 have been tested with Wide-ResNet-28-10, Shake-Shake 26 2x32d and PyramidNet backbones; SVHN with Wide-ResNet-28-10 backbone; and ImageNet with ResNet-50 and ResNet-200 backbones.
For the sake of clarity, we computed and compared the gain in accuracy for all the methods based on quantitative results reported in the various original works. Overall, the results show very high performance for automated data augmentation methods over classical methods. In Figure 15, we compare the performance of the best automated data augmentation methods with state-of-the-art approaches based on classical techniques on CIFAR-10. In the case of classical methods, IHDA [195] shows the highest performance (+1.7%), followed by MixUP (+1.2%) and GridMask (+1.1%). These are all well-below the best automated augmentation methods, which achieve an average of +2.17% gain in classification accuracy. In Figure 15, it can be observed that only one classical data augmentation method, IHDA [195] at 6, ranks among the top ten data augmentation methods. Similarly, for CIFAR-100 (Figure 16), the automated methods have convincingly outperformed the classical augmentation methods, averaging +1.47% gain in performance compared with +2.28% for traditional methods. Only two classical methods–SuperMix(+2.4%) and GridMask (+2.2%)–are part of the ten best performing augmentation methods. Figure 17 shows the relative performance of automated and trational methods on ImageNet dataset with ResNet-50 backbone. Results are shown for top-1% and top-5% accuracies. Here, too, AutoMl-based augmentation methods significantly outperform their classical counterparts. For instance, For the top-5% classification accuracies, four out of the top five and eight out of the top ten methods are automated augmentation methods. These results convincingly demonstrate the superior performance of data augmentation techniques based AutoML pipelines.
@@ 7.3.3 Improving performance by combining classical and AutoML methods
They first assessed the performance of each technique separately before pairing AutoAugment with each of the classical augmentation methods. The results obtained by their experiments showed marked improvement by the paired augmentation strategies over AutoAugment. The combined augmentations also outperformed each of the classical methods when applied independently. In all cases of combined augmentations resulted in significant performance improvements over corresponding single augmentations. More importantly, utilizing a combination of three augmentations also showed a marked improvement over those employing only two augmentation strategies. The results of the study are summarized in Table VI. The author used CIFAR10 dataset with WideResNet-40-2 backbone in their experimental setting. Note that the original works AA, MixUp and CutMix do not report results for the WideResNet-28-10 backbone. Because of the difference in training settings, the accuracies of individual methods reported (and shown in the Figure 15) are slightly different from those obtained by the original sources. Nonetheless, they reflect the general performance of the models based on the specific techniques and dataset configurations. The experimental results show that while automated data augmentation methods hold significant promise, there is still scope for the application of classical methods to complement augmentations generated automatically. However, applying multiple augmentations in this manner ought to be carefully considered as combining augmentations have been shown to harm performance in some settings (e.g., [200]).
@@ 8.3 Open problems and research directions
  • One of the most challenging tasks in the implementation of automated data augmentation, and indeed in the development of AutoML models in general, is the task of composing a good search space that can provide effective solution for the target application. Even though approaches that learn to automatically generate effective search spaces from data have been devised, they generally require substantial domain knowledge to design, making it challenging for nonexperts to accomplish. Recently, large language models (LMMs) have been leveraged to automate complex machine learning tasks. Preliminary work by Yang et al. [204] has shown that with LLMs, it will be possible to fully automate the development of AutoML and operation of frameworks in a way that will allows lay users to build and use these models. For instance, leveraging the rich knowledge of LLMs will allow nonexpert developers to easily generate effective, context-relevant search spaces and associated hyperparameters for data augmentation. In addition, intuitive user interfaces based on LLMs will enable lay users to simply specify natural language instructions to configure various settings of the resulting model. These frameworks, through the LMM interfaces, can also provide useful feedback and suggestions, as well as explanations about the internal mechanisms and operation of the model.
@@
  • The automation of data augmentation is an extremely computationally intensive process. Given that for a dataset several augmentations can be combined in different ways (e.g., varying numbers of transformation operations and different ordering of their application) and each transformation operation can be applied with infinitely wide range of intensities, the number of possible augmentations is effectively unlimited. In order to establish the effectiveness of any new augmentation using appropriate search strategies, it is usually necessary to conduct exhaustive training and subsequently validate the performance on a proxy task or the target task. This procedure is a combinatorial problem and, using current approaches, in some cases it is impractical to arrive at an optimal solution within a practical time frame. A potentially effective workaround is to leverage large language model prompts to guide the search process towards better solutions. Such a mechanism could also provide a means of high-level interaction that allows developers’ priorities and constraints (e.g., infrastructure limitations, economic constraints, time laxity, etc.) to be factored in the search process. The performance evaluation step could also benefit from such new techniques by leveraging real-time feedback from
@@
  • One of the major problems of machine learning tasks that require data augmentation is imbalanced data—a situation where some classes underrepresented while others are overrepresented. This results in a situation where predictive performance on minority classes is severely compromised. Even though there are data augmentation-based workarounds specifically designed to overcome this problem (i.e., to balance imbalanced data), automated data augmentation methods have not been extended to this domain. Future research is expected to produce dedicated AutoML frameworks specifically aimed at balancing imbalance training data.
@@
  • While conventional data augmentation strategies rely on constructing dataset- and task-specific solutions, automatically learned policies have been shown (e.g., [17] to be generally more transferable to new datasets. This property could potentially be exploited to provide further insights towards developing generic augmentations strategies for a wider range of deep learning tasks that are independent of datasets. Some interesting results [205] in this direction have already been reported. Future line of research will involve devising mechanisms to optimize models based on their transferability. This will enable the underlying models to encode more transferable architectureual features and hyperparameter settings, thereby generating task- and dataset-agnostic AutoML frameworks.

### v3

@@ 5.6 Searchless optimization of data augmentation policies
Although the optimization techniques discussed in Subsections 5.1 through 5.4 have demonstrated impressive results, they are still too expensive for most practical purposes. For instance, optimization methods employed in AutoAugment takes about 15,000 GPU hours to train on ImageNet [131]. To address the challenge of computational complexity, some recent works (e.g., [34, 170, 171, 162]) propose simplified search space design where good augmentation policies can be found using much simpler search strategies like random search [172], grid search [173] or greedy search [174]. Greedy AutoAugment [170] proposes to reduce the search space by eliminating the need to search for different combinations of augmentation operations in a combinatorial manner. The basic idea is to dynamically expand the space of augmentation operations in the direction of effective augmentation policies by creating and modifying sub-policies with good performance. In the search process, only transformation type and intensity are the only hyperparameteres considered. The probability of applying the operations are determined later when the search is complete and is only applied on the effective policies found. This significantly reduces the amount of computations needed to find good augmentation strategies. Because of its simplicity and effectiveness, the greedy search policy optimization procedure has been used in several works [175].
@@
Unlike in RA [34] and all previous optimization strategy formulations, Uniform Augment (UA) [162] completely bypasses the search stage and instead propose to uniformly sample augmentation policy hyperparameters from the search space. To achieve this, the authors formulate the search space as distribution invariant by ensuring that the transformation operations and associated intensities result in label-preserving augmentations, and the transformed data remains within the distribution of the original input data. A machine learning model is then able to learn optimal augmentations by sampling randomly from the search space and optimizing model parameters using gradient descent. Uniform Augment, thus far, is the cheapest in terms of the number of computations needed to find optimal augmentation policies. Indeed, the number of search operations is theoretically zero.
@@ 6 Evaluation methods
An important step after selecting a subset of operations for the search space and an optimization method is to evaluate the performance of the resulting models. This process is repeated for different combinations of model configuration settings (i.e., all plausible sets of augmentation operations, hyperparameters and search methods). At the end of the process, performance results are compared and an optimal pipeline and hyperparameters chosen according to a specified performance criterion. The costly nature of the evaluation stage significantly limits the range of plausible configurations that can be explored to achieve optimal predictive performance. Instead of training the network to converge before evaluating its performance, approaches have been devised to accelerate the process by enabling near-optimal strategies to be found without exhaustive evaluations.
One of the most popular approaches is the so-called early stopping strategy [23], a technique based on terminating the evaluation process for settings that are predicted to perform relatively poor on a validation set. Some works, for instance, RandAugment (RA) [34], TrivialAugment (TA) [171] and OHLAA [114], reduce the range of augmentation operations and hyperparameters in the search space in the evaluation process. A popular method to speed up model evaluation is to reduce the fidelity of the input data [23]. In image augmentation, this is achieved by reducing the sizes or resolutions of the images used. Another common approach is to train on proxy tasks–i.e., employ reduced model size or subsets of training data. Examples of AutoML-based data augmentation models that employ this strategy include AutoAugment (AA) [17], FastAA [108], [108], FasterAA [30], DADA [159], AWS [145], and PBA [37]. Another method for reducing the enormous computaional overhead at the evaluation stage is to employ surrogate models as evaluators [176, 177] to predict the performance of the target models. This approach circumvents the need to perform costly evaluation on different configurations of the real model.
@@ 7 Quantitative performance of data augmentation approaches
We present results on the performance of AutoML-based data augmentation methods and compare these with state-of-the-art data augmentation methods based on classical approaches. For fairness we compare methods that use the same datasets and similar training settings (i.e., similar model configuration and number of training epochs). The results presented here are curated from the original works. We highlight instances where results or other performance information from secondary sources are used.
@@ 7.1.1 Datasets
The CIFAR-10 [111] and CIFAR-100 [111] are highly popular image classification datasets. Each of the datasets contains 60,000 labeled 32x32 RGB images that are divided into 50,000 training and 10,000 test images. The images in CIFAR-10 are divided into 10 classes of 6,000 images per class, while CIFAR-100 has 100 classes, each containing 600 (500 training and 10 test) images.
The Street View House Numbers (SVHN) [112] is a digit (i.e., decimal numbers 0 to 9) recognition dataset containing 600,000 color (RGB) images. The dataset is made up of cropped, 32×32 images derived from Google Street View [180]. Thus, the images are pictures of real-world house number plates in challenging settings. SVHN include bounding box information in a separate file and can therefore be used for digit detection in addition to image classification.
@@
The Microsoft Common Objects in Context (MS COCO) dataset [179] is a large-scale dataset that can be used for many computer vision tasks. It is primarily designed for object detection, pose estimation, keypoint detection, image captioning, and segmentation tasks. It has a total of 330,000 images split into 80 classes.
@@ 7.1.2 Performance metrics and model settings
For the classification tasks, the performance metric commonly used is the generalization accuracy. Results for ImageNet include top-1% and top-5% accuracy measures. Mean average precision (mAP) is used to characterize performance of the object detection tasks, i.e., on the MS COCO dataset (Table V).
Performance on CIFAR-10 and CIFAR-100 has been tested using three different backbone models: Wide Residual Networks [182], specifically, Wide-ResNet-28-10 (see the original work for a detailed description of this setting); Shake-Shake (26 2x32d) [183]; and PyramidNet [184]. Results for ImageNet dataset have are based on ResNet-50 and ResNet-200 backbone models (see He et al. [185] for details). RetinaNet [186] is used with various ResNet models [187] on the MS COCO dataset.
@@ 7.2 Performance of AutoML-based data augmentation methods
Most of the existing works report results for image classification using large-scale CNN models. The most popular datasets used for these purposes are the ImageNet and CIFAR (CIFAR-10 and CIFAR-100) family of datasets described in Subsection 7.1. Detailed performance results for automated data augmentation methods based on these datasets are summarized in Tables Tables III, IV and VII. Table III show image classification results for CIFAR-10 and CIFAR-10 for three different backbone models– WideResNet 28-10 (WRN28-10) [182], Shake-Shake 26-32 (S-S 26-32) [183] and PyramidNet (PNet) [184]. For each data augmentation method, we indicate the average classification accuracy for the three backbone models. From the Table, the mean gain in accuracy range between 0.6 and 1.5 on CIFAR-10 and between 0.8% and 3.1% on CIFAR-100. This demonstrates consistentently strong performance for all methods. Indeed, the average performance gain for all the methods is 1.02% and 2.17% on CIFAR-10 and CIFAR-100, respectively. Note that for CIFAR-100, many of the methods do not have results for PyramidNet backbone on CIFAR-100. We exclude these methods in our average performance computations. Similar to Table III, Table IV shows classification results for the ImageNet dataset based on ResNet-50 and ResNet-200 backbone models. Results are shown for top-1 and top-5% accuracies. The table also indicates percentage performance gain for each of these accuracy measures. The gains are consistent for all methods, with the lowest top-1% accuracy gain on RestNet-50 being +0.9% (for PBA method) and +1.5% on ResNet-200. The performance gains for various AutoML-based data augmentation methods on the SVHN dataset are summarized in Figure 14. Again, the results show impressive performance on challenging digit classification tasks. In Table VII we compute the average performance of different categories of automated data augmentation methods based on the optimization strategies they employ. The results show that performance difference across different search methods, on average, is largely insignificant.
While the majority of approaches have reported results for image classification tasks, a few works have evaluated the performance of automated data augmentation techniques on machine learning tasks other than image classification. For instance, Liu et al. [107], Cubuk et al. [34] and Chen et al. [188] investigate the performance of common automated data augmentation methods–AA [17], DADA [31], DDAS [107]–on object detection tasks. The tests were conducted using large-scale detection models, specifically, ResNet-101 and RetinaNet. The results of these experiments are summarized in Table V. These studies show that AutoML-based data augmentation methods are effective for object detection. Indeed, per the results in Table V, the data augmentation techniques lead to an average performance improvement of 1.54% (mean average precision or mAP) for object detection models.
@@ 7.3.2 Comparison of automated data augmentation methods and state-of-the art classical approaches
In this section, we compare the performance of state-of-the-art data agumentation methods based on classical approaches with automated data augmentation methods. The results are summarized in Figures 15, 16 and 17. Again, to ensure a uniform point of reference for the assessment, we selected methods that have been evaluated on the same datasets and CNN backbone models for comparison. Specifically, the results provided for CIFAR-10 and CIFAR-100 have been tested with Wide-ResNet-28-10, Shake-Shake 26 2x32d and PyramidNet backbones; SVHN with Wide-ResNet-28-10 backbone; and ImageNet with ResNet-50 and ResNet-200 backbones.
For the sake of clarity, we computed and compared the gain in accuracy for all the methods based on quantitative results reported in the various original works. Overall, the results show consistently high performance for automated data augmentation methods over classical methods. In Figure 15, we compare the performance of the best automated data augmentation methods with state-of-the-art approaches based on classical techniques on CIFAR-10. In the case of classical methods, IHDA [198] shows the highest performance (+1.7%), followed by MixUP (+1.2%) and GridMask (+1.1%). These are all well-below the best automated augmentation methods, which achieve an average of +2.17% gain in classification accuracy. In Figure 15, it can be observed that only one classical data augmentation method, IHDA [198] at
${6^{th}}$
, ranks among the top ten data augmentation methods. Similarly, for CIFAR-100 (Figure 16), the automated methods have convincingly outperformed the classical augmentation methods, averaging +1.47% gain in performance compared with +2.28% for classical methods. Only two classical methods–SuperMix(+2.4%) and GridMask (+2.2%)–are part of the ten best performing augmentation methods. Figure 17 shows the relative performance of automated and trational methods on ImageNet dataset with ResNet-50 backbone. Results are shown for top-1% and top-5% accuracies. Here, too, AutoMl-based augmentation methods significantly outperform their classical counterparts. For instance, for the top-5% classification accuracies, four out of the top five and eight out of the top ten methods are automated augmentation methods. These results convincingly demonstrate the superior performance of data augmentation techniques based AutoML pipelines.
@@ 7.3.3 Improving performance by combining classical and AutoML methods
They first assessed the performance of each technique separately before pairing AutoAugment with each of the classical augmentation methods. The results obtained by their experiments showed marked improvement by the paired augmentation strategies over AutoAugment. The combined augmentations also outperformed each of the classical methods when applied independently. All cases of combined augmentations resulted in significant performance improvements over corresponding single augmentations. More importantly, utilizing a combination of three augmentations also showed a marked improvement over those employing only two augmentation strategies. The results of the study are summarized in Table VI. The author used CIFAR10 dataset with WideResNet-40-2 backbone in their experimental setting. Note that the original works AA, MixUp and CutMix do not report results for the WideResNet-28-10 backbone. Because of the difference in training settings, the accuracies of individual methods reported (and shown in the Figure 15) are slightly different from those obtained by the original sources. Nonetheless, they reflect the general performance of the models based on the specific techniques and dataset configurations. The experimental results show that while automated data augmentation methods hold significant promise, there is still scope for the application of classical methods to complement augmentations generated automatically. However, applying multiple augmentations in this manner ought to be carefully considered as combining augmentations have been shown to harm performance in some settings (e.g., [203]).
@@ 8.3 Open problems and research directions
  • One of the most challenging tasks in the implementation of automated data augmentation, and indeed in the development of AutoML models in general, is the task of composing a good search space that can provide effective solution for the target application. Even though approaches that learn to automatically generate effective search spaces from data have been devised, they generally require substantial domain knowledge to design, making it challenging for nonexperts to accomplish. Recently, large language models (LMMs) have been leveraged to automate complex machine learning tasks. Preliminary work by Yang et al. [207] has shown that with LLMs it will be possible to fully automate the development of AutoML and operation of frameworks in a way that allows lay users to build and use these models. For instance, leveraging the rich knowledge of LLMs will allow nonexpert developers to easily generate effective, context-relevant search spaces and associated hyperparameters for data augmentation. In addition, intuitive user interfaces based on LLMs will enable lay users to simply specify natural language instructions to configure various settings of the resulting model. These frameworks, through the LMM interfaces, can also provide useful feedback and suggestions, as well as explanations about the internal mechanisms and operation of the model.
@@
  • The automation of data augmentation is an extremely computationally intensive process. Given that for a dataset several augmentations can be combined in different ways (e.g., varying numbers of transformation operations and different ordering of their application) and each transformation operation can be applied with infinitely wide range of intensities, the number of possible augmentations is effectively unlimited. In order to establish the effectiveness of any new augmentation using appropriate search strategies, it is usually necessary to conduct exhaustive training and subsequently validate the performance on a proxy task or the target task. This procedure is a combinatorial problem and, using current approaches, in some cases it is impractical to arrive at an optimal solution within a practical time frame. A potentially effective workaround is to leverage large language model prompts to guide the search process towards better solutions. Such a mechanism could also provide a means of high-level interaction that allows developers’ priorities and constraints (e.g., infrastructure limitations, economic constraints, time laxity, etc.) to be factored in the search process. The performance evaluation step could also benefit from such new techniques by leveraging real-time feedback from LLMs.
@@
  • One of the major problems of machine learning tasks that require data augmentation is imbalanced data—a situation where some classes are underrepresented while others are overrepresented. This results in a problem where predictive performance on minority classes is severely compromised. Even though there are data augmentation-based workarounds specifically designed to overcome this problem (i.e., to balance imbalanced data), automated data augmentation methods have not yet been extended to this domain. Future research is expected to produce dedicated AutoML frameworks specifically aimed at balancing imbalance training data.
@@
  • While conventional data augmentation strategies rely on constructing dataset- and task-specific solutions, automatically learned policies have been shown (e.g., [17]) to be generally more transferable to new datasets. This property could potentially be exploited to provide further insights towards developing generic augmentations strategies for a wider range of deep learning tasks that are independent of datasets. Some interesting results [208] in this direction have already been reported. Future line of research will involve devising mechanisms to optimize models based on their transferability. This will enable the underlying models to encode more transferable architectureual features and hyperparameter settings, thereby generating task- and dataset-agnostic AutoML frameworks.